# Towards Adapting Federated & Quantum Machine Learning for Network Intrusion Detection: A Survey

### v1

@@ V-C3 Quantum Reinforcement Learning (QRL)
As in classical reinforcement learning, in QRL, the agent learns to make decisions by interacting with an environment to maximize cumulative reward. Like traditional reinforcement learning (RL), a QRL is built with three components: a policy, a reward function, and a model of the environment [138]. However, QRL differs from RL in how it represents states and actions using quantum states in high-dimensional Hilbert space. Each action or state can exist in a superposition of multiple eigen states allowing exploration of many possibilities simultaneously. Action in QRL is based on quantum measurement and values across all states can be updated simultaneously utilizing quantum parallelism. The learning process involves updating the probability amplitudes of actions based on received rewards. In IDS, where network traffic data is complex and high-dimensional, traditional RL methods may struggle with scalability and precision. QRL utilizing quantum parallelism can be highly scalable and efficient in learning high-dimensional features and intricate patterns of network traffic data.
@@ V-D Case Studies of QML in IDS
Abreu et al. [139] UNSW-NB15, CIC-IDS17, CICIoT23, TON IoT VQC, QKM, QCNN, QSVM Qiskit F1 score
Kumar et al. [140] EDGE-IIOTSET, ACI IoT QSVC Qiskit Recall, Precision, F1 score
Hdaib et al. [104] KDD99, IoT-23, CIC-IoT-23 QAE with QSVM, QAE with Quantum Random Forest, QAE with QkNN Qiskit, PennyLane PyTorch Accuracy, Precision, Recall, F1 score
Rahman et al. [141] NSL-KDD qGAN Qiskit, PyTorch Cumulative Distribution Function (CDF)
@@
In [104], Hdaib et al. proposed three different quantum autoencoder (QAE) frameworks to detect anomalies in network traffic. The three proposed frameworks are formed by integrating QAE with a quantum one-class support vector machine, a quantum random forest and a quantum $k$-nearest neighbor, and are tested on three different benchmark datasets comprising network flows. A QAE can be trained on ”normal” data samples to reconstruct the original data from the compressed state. The primary goal is to minimize the loss of reconstruction. The high reconstruction loss shows an anomaly in the network flow. The QAE alone achieved an accuracy of 75% and an F1 score of 77%. However, when integrated with other QML methods, it significantly outperformed the baseline, demonstrating high detection capabilities and robustness. Among the three proposed frameworks, QAE when combined with QkNN achieved the best performance, with an accuracy of 97.79%, and an F1 score of 98.26%. The authors concluded that quantum deep learning when integrated with parameterized quantum circuits can significantly enhance anomaly detection performance.
@@
In [140], Kumar et al. developed an IDS for IoT networks based on a quantum support vector classifier (QSVC) namely QuIDS. Two publicly available IoT network traffic datasets were used. The architecture of the proposed module included three main modules: Flow Reconstruction, Feature Extraction, and QSVC Classification. The input to the flow reconstruction module was a set of packets $P=\{P_{1},P_{2},P_{3},...,P_{n}\}$ and the output was the set of flows $F=\{F_{1},F_{2},F_{3},...,F_{m}\}$ were $n>m$. The set of flows, F, is given as input to the feature extraction module which processes these flows and outputs flow-level features vector given by $X=\{\vec{x_{1}},\vec{x_{2}},\vec{x_{3}},...,\vec{x_{8}}\}$. This module extracts eight flow-level generic traffic features such as Flow Direction Ratio (FDiR), Flow Payload Ratio (FPR), Flow Duration Ratio (FDuR), Flow Packet Rate Ratio (FPRR), Flow Inter-arrival Time Ratio (FITR), Flow Packet Length Ratio (FPLR), Flow Length Rate Ratio (FLRR), and Flow TLS Record Length Ratio (FTRLR) that is to be utilized for attack classification in IoT network traffic. In the QSVC classification module, features are reduced to optimal qubits using PCA. The features are mapped to quantum states using three different feature mapping techniques (i.e., ZFeature, ZZFeature and PauliFeature maps). After encoding features to quantum states, the QSVC was trained using the COBYLA optimizer. The results were compared with five machine learning, three deep learning and two recent state-of-the-art methods. The results showed that QuIDS using the ZZFeature map outperformed all the other compared models in terms of precision, recall and F1-score.
@@
Rahman et al. in [141] implemented a quantum generative adversarial network (qGAN) using PyTorch in conjunction with Qiskit. In the proposed architecture, the generator was modeled as a quantum circuit and the discriminator remained classical. Binary cross-entropy was used as a loss function which is given by:
@@
In [139], Abreu et al. proposed QuantumNetSec, a QML-based network security framework specifically tailored for Noisy Intermediate-Scale Quantum (NISQ) devices. The QuantumNetSec algorithm begins with monitoring real-time network traffic and metadata collection. The raw data is preprocessed using classical methods and mapped to quantum states using a feature map. It is followed by transpilation and circuit optimization techniques adapted to the NISQ backend. The proposed algorithm was tested using four different QML models: variational quantum circuits (VQC), quantum k-means (QKM), quantum convolutional neural networks (QCNN), and quantum support vector machines (QSVM). These quantum models were optimized using classical optimizers. In the proposed method four feature maps (RawFeatureVector, PauliFeatureMap, ZFeatureMap and ZZFeatureMap), four Ansatz (RealAmplitudes, EfficientSU2, ExcitationPreserving and TwoLocal), and four different optimizers (ADAM, COBYLA, SPSA and gradient descent) were employed. A key innovation in this method is the personalized optimization technique which tests all the combinations of transpilation methods, such as init, layout, translation, routing, optimization and scheduling, to reduce the noise in the quantum hardware. For evaluation, seven different backends from the IBM quantum platform were used: Cairo, Kyoto, Brisbane, Osaka, Sherbrooke, Torino, and Quebec. The results showed that the personalized optimization technique increased the performance of all four QML models, achieving a higher F1 score, and QCNN achieved the highest F1 score in both binary and multiclass classification.
@@ VI Transitioning from QML to QFL
Qiskit [142] IBM Python Yes Yes Terra, Aer, Aqua; supports hybrid QML models and integration with FL via PySyft or Flower.
Cirq [143] Google Python Yes Possible Low-level gate model; integrated with TensorFlow Quantum; flexible for hybrid ML.
Strawberry Fields [144] Xanadu Python Yes Experimental (via PennyLane) Photonic (CV) quantum computing; used with PennyLane for hybrid QML.
QuTiP [145] Open-source Python Limited No Simulates open quantum systems; not designed for ML or FL.
PennyLane [146] Xanadu Python Strong Yes (via Flower, PySyft) Hybrid QML; supports Braket, Qiskit, Cirq; used in QFL research with Flower/PySyft.
TF Quantum [147] Google Python Yes Possible (via TensorFlow Federated) TensorFlow + Cirq for hybrid models; extendable to QFL via TensorFlow Federated.
MS QDK [148] Microsoft Q# Basic No Designed for algorithm prototyping in Q#; lacks ML/FL support.
Amazon Braket [149] AWS Python Yes Yes Access to IonQ, OQC, Rigetti; PennyLane integration; QFL possible via custom FL orchestration.
@@ VI-A1 QFL for Cross-Silo Network Security
Cross-Silo QFL lets various organizations (e.g., banks, government agencies or ISPs) collaboratively detect intrusions or attacks without sharing raw network traffic. Each organization has substantial infrastructure (HPC clusters, quantum simulators, or priority access to cloud QPUs). Each silo collects network traffic from its own perimeter. These classical network traffic data are mapped to quantum states using feature mapping techniques, and a quantum model is trained locally across each organization. Only quantum circuit parameters or model weights are securely aggregated at a central coordinator. Techniques such as QKD [150] and quantum secure multi-party computation [151] can be incorporated to ensure the security of the shared-parameters while communicating with the server.
@@ VI-B Communication in QFL
Communication in QFL can be purely classical (with quantum local training), fully quantum (with quantum state sharing), or hybrid. The choice depends on hardware constraints, desired security, and application domain. In the near term, classical communication secured with QKD [152, 153] is most practical, whereas entanglement-assisted quantum communication represents the long-term vision.
@@
To explore these modalities in more depth, we distinguish between classical communication and quantum communication in QFL [154].
@@ VI-C Secure Parameter Sharing in QFL
An efficient and secure communication is one of the major challenges of FL. Classical encryption techniques provide a certain level of protection, but they may fail with increasingly sophisticated attacks. Quantum-secure multi-party computation (QSMPC) [151] approaches in QFL play a foundational role by enabling multiple parties to jointly compute model updates without revealing their private data. Building upon this, protocols such as Quantum Key Distribution (QKD) [150] and Quantum Secure Direct Communication (QSDC) [155] leverage quantum mechanics principles, like the irreversibility of measurement and photon polarization, to safeguard against eavesdropping by detecting interception through elevated quantum bit error rates. Additionally, quantum teleportation [156] allows the transfer of quantum states without physically transmitting the qubits themselves. Using pre-shared entanglement combined with classical communication, teleportation facilitates highly secure quantum parameter sharing, especially over long or noisy channels. Finally, quantum techniques such as quantum error correction (QEC) [157] can be utilized to preserve the integrity and reliability of quantum data exchanged via teleportation, QKD, or QSDC. Various quantum techniques can be used to compress the model to reduce the communication overhead.
@@ VI-D1 FedQNN
The FedQNN framework proposed by Innan et al. in [158] extends federated learning by integrating quantum neural networks (QNNs) to enable collaborative model training while maintaining local data privacy. In this approach, each client retains its private dataset and encodes it into quantum states processed through PQCs. The quantum model updates, represented as optimized quantum parameters, are communicated to a central server rather than the raw data. The server aggregates these updates to form a global quantum model, mathematically expressed as the average of client parameters, which enhances overall predictive performance and accuracy. Within the QNN, Hadamard gates create superposition, rotation gates introduce phase shifts and superpositions along X and Y axes, and the $U(\theta)$blocks iteratively optimize network parameters. Measurements on the final quantum state produce predictions, which are evaluated using a mean squared error (MSE) loss function. FedQNN also incorporates secure communication protocols and privacy-preserving aggregation mechanisms, allowing clients to contribute to the global model without exposing sensitive data. This integration of QNNs with federated learning leverages the computational advantages of quantum circuits while addressing the challenges of distributed and privacy-sensitive machine learning environments.
@@ VI-D2 Noise Aware Clustered Quantum Federated Learning (NAC-QFL)
In [159], Sahu et al. proposed NAC-QFL. The NAC-QFL system integrates a QML model with a parameterized quantum circuit into a federated learning framework across a set of quantum devices organized into clusters to optimize communication efficiency. Each cluster has a cluster head responsible for device selection, local model training, and coordination with a central aggregation server. Clustering uses a K-means algorithm with a distance metric combining inter-device communication distance and entanglement-assisted channel capacity, and cluster heads are selected based on both channel capacity and classical computational resources. The selection of devices within clusters is sensitive to noise, and the effective noise of each device is modeled considering qubit relaxation, dephasing, gate errors, measurement errors, and state preparation errors. The least noisy devices that satisfy circuit capacity and parallelization constraints are selected for training. During NAC-QFL training, the global model is distributed from the server to cluster heads, propagated to selected devices for local training, and aggregated back at the server, ensuring scalable, robust, and noise-resilient federated quantum learning.
@@ VI-D3 Multimodal Quantum Federated Learning Framework with Fully Homomorphic Encryption (MQFL-FHE)
Dutta et al. in [160] proposed MQFL-FHE, which enables secure federated learning across heterogeneous data modalities, such as images, text, or biomedical sequences, while preserving client privacy. In this framework, each client preprocesses its local multimodal data and encodes it into quantum states using parameterized unitary operations. Local quantum model updates are computed via PQCs and subsequently encrypted using the CKKS fully homomorphic encryption (FHE) scheme, allowing the central server to aggregate encrypted model updates without accessing the raw data. Feature-level fusion is performed locally using attention-based mechanisms to capture correlations across modalities, and the aggregated global model is updated homomorphically by weighting client contributions according to their dataset size. This integration of quantum computation and FHE mitigates the performance degradation typically introduced by encryption while ensuring that sensitive data remains private during the federated learning process. The approach is scalable to multiple modalities and allows each client to contribute securely to a global model that captures multimodal relationships effectively.
@@ VI-D5 Chained Continuous Quantum Federated Learning (ccQFL)
In [162], Gurung et al. proposed ccQFL. In QFL, a central server is responsible for performing aggregation of parameters to produce a global model making it a single point of failure if the server is compromised. The constant communication between a central server and clients may not be feasible in the real world. ccQFL addresses this issue by performing serverless aggregation. Authors have introduced two variants of ccQFL. In the first variant, the devices are randomly grouped or selected based on some selection criteria. The devices train the models in parallel within the group. After the parallel training within the group is completed, the intra-group training is conducted sequentially without the need for an aggregation mechanism. In the second variant, $n$ devices are grouped in pairs. The model is first trained on the first device of each pair and passed on to the second device of the same pair for further training. The trained model from the second device is passed to the untrained second device in alternate pairs. Finally, the models of these devices are averaged and the process continues until the training is completed. This method ensures serverless aggregation of the model and also helps in dealing with heterogeneity and non-IID data.

### v2

@@ V-C3 Quantum Reinforcement Learning (QRL)
As in classical reinforcement learning, in QRL, the agent learns to make decisions by interacting with an environment to maximize cumulative reward. Like traditional reinforcement learning (RL), a QRL is built with three components: a policy, a reward function, and a model of the environment [dong2008quantum]. However, QRL differs from RL in how it represents states and actions using quantum states in high-dimensional Hilbert space. Each action or state can exist in a superposition of multiple eigen states allowing exploration of many possibilities simultaneously. Action in QRL is based on quantum measurement and values across all states can be updated simultaneously utilizing quantum parallelism. The learning process involves updating the probability amplitudes of actions based on received rewards. In IDS, where network traffic data is complex and high-dimensional, traditional RL methods may struggle with scalability and precision. QRL utilizing quantum parallelism can be highly scalable and efficient in learning high-dimensional features and intricate patterns of network traffic data.
@@ V-D Case Studies of QML in IDS
Abreu et al. [abreu2025quantumnetsec] UNSW-NB15, CIC-IDS17, CICIoT23, TON IoT VQC, QKM, QCNN, QSVM Qiskit F1 score
Kumar et al. [kumar2025quids] EDGE-IIOTSET, ACI IoT QSVC Qiskit Recall, Precision, F1 score
Hdaib et al. [hdaib2024quantum] KDD99, IoT-23, CIC-IoT-23 QAE with QSVM, QAE with Quantum Random Forest, QAE with QkNN Qiskit, PennyLane PyTorch Accuracy, Precision, Recall, F1 score
Rahman et al. [rahman2023quantum] NSL-KDD qGAN Qiskit, PyTorch Cumulative Distribution Function (CDF)
@@
In [hdaib2024quantum], Hdaib et al. proposed three different quantum autoencoder (QAE) frameworks to detect anomalies in network traffic. The three proposed frameworks are formed by integrating QAE with a quantum one-class support vector machine, a quantum random forest and a quantum $k$-nearest neighbor, and are tested on three different benchmark datasets comprising network flows. A QAE can be trained on ”normal” data samples to reconstruct the original data from the compressed state. The primary goal is to minimize the loss of reconstruction. The high reconstruction loss shows an anomaly in the network flow. The QAE alone achieved an accuracy of 75% and an F1 score of 77%. However, when integrated with other QML methods, it significantly outperformed the baseline, demonstrating high detection capabilities and robustness. Among the three proposed frameworks, QAE when combined with QkNN achieved the best performance, with an accuracy of 97.79%, and an F1 score of 98.26%. The authors concluded that quantum deep learning when integrated with parameterized quantum circuits can significantly enhance anomaly detection performance.
@@
In [kumar2025quids], Kumar et al. developed an IDS for IoT networks based on a quantum support vector classifier (QSVC) namely QuIDS. Two publicly available IoT network traffic datasets were used. The architecture of the proposed module included three main modules: Flow Reconstruction, Feature Extraction, and QSVC Classification. The input to the flow reconstruction module was a set of packets $P=\{P_{1},P_{2},P_{3},...,P_{n}\}$ and the output was the set of flows $F=\{F_{1},F_{2},F_{3},...,F_{m}\}$ were $n>m$. The set of flows, F, is given as input to the feature extraction module which processes these flows and outputs flow-level features vector given by $X=\{\vec{x_{1}},\vec{x_{2}},\vec{x_{3}},...,\vec{x_{8}}\}$. This module extracts eight flow-level generic traffic features such as Flow Direction Ratio (FDiR), Flow Payload Ratio (FPR), Flow Duration Ratio (FDuR), Flow Packet Rate Ratio (FPRR), Flow Inter-arrival Time Ratio (FITR), Flow Packet Length Ratio (FPLR), Flow Length Rate Ratio (FLRR), and Flow TLS Record Length Ratio (FTRLR) that is to be utilized for attack classification in IoT network traffic. In the QSVC classification module, features are reduced to optimal qubits using PCA. The features are mapped to quantum states using three different feature mapping techniques (i.e., ZFeature, ZZFeature and PauliFeature maps). After encoding features to quantum states, the QSVC was trained using the COBYLA optimizer. The results were compared with five machine learning, three deep learning and two recent state-of-the-art methods. The results showed that QuIDS using the ZZFeature map outperformed all the other compared models in terms of precision, recall and F1-score.
@@
Rahman et al. in [rahman2023quantum] implemented a quantum generative adversarial network (qGAN) using PyTorch in conjunction with Qiskit. In the proposed architecture, the generator was modeled as a quantum circuit and the discriminator remained classical. Binary cross-entropy was used as a loss function which is given by:
@@
In [abreu2025quantumnetsec], Abreu et al. proposed QuantumNetSec, a QML-based network security framework specifically tailored for Noisy Intermediate-Scale Quantum (NISQ) devices. The QuantumNetSec algorithm begins with monitoring real-time network traffic and metadata collection. The raw data is preprocessed using classical methods and mapped to quantum states using a feature map. It is followed by transpilation and circuit optimization techniques adapted to the NISQ backend. The proposed algorithm was tested using four different QML models: variational quantum circuits (VQC), quantum k-means (QKM), quantum convolutional neural networks (QCNN), and quantum support vector machines (QSVM). These quantum models were optimized using classical optimizers. In the proposed method four feature maps (RawFeatureVector, PauliFeatureMap, ZFeatureMap and ZZFeatureMap), four Ansatz (RealAmplitudes, EfficientSU2, ExcitationPreserving and TwoLocal), and four different optimizers (ADAM, COBYLA, SPSA and gradient descent) were employed. A key innovation in this method is the personalized optimization technique which tests all the combinations of transpilation methods, such as init, layout, translation, routing, optimization and scheduling, to reduce the noise in the quantum hardware. For evaluation, seven different backends from the IBM quantum platform were used: Cairo, Kyoto, Brisbane, Osaka, Sherbrooke, Torino, and Quebec. The results showed that the personalized optimization technique increased the performance of all four QML models, achieving a higher F1 score, and QCNN achieved the highest F1 score in both binary and multiclass classification.
@@ VI Transitioning from QML to QFL
Qiskit [cross2018ibm] IBM Python Yes Yes Terra, Aer, Aqua; supports hybrid QML models and integration with FL via PySyft or Flower.
Cirq [omole2020cirq] Google Python Yes Possible Low-level gate model; integrated with TensorFlow Quantum; flexible for hybrid ML.
Strawberry Fields [killoran2019strawberry] Xanadu Python Yes Experimental (via PennyLane) Photonic (CV) quantum computing; used with PennyLane for hybrid QML.
QuTiP [johansson2012qutip] Open-source Python Limited No Simulates open quantum systems; not designed for ML or FL.
PennyLane [bergholm2018pennylane] Xanadu Python Strong Yes (via Flower, PySyft) Hybrid QML; supports Braket, Qiskit, Cirq; used in QFL research with Flower/PySyft.
TF Quantum [broughton2020tensorflow] Google Python Yes Possible (via TensorFlow Federated) TensorFlow + Cirq for hybrid models; extendable to QFL via TensorFlow Federated.
MS QDK [QDK] Microsoft Q# Basic No Designed for algorithm prototyping in Q#; lacks ML/FL support.
Amazon Braket [braket] AWS Python Yes Yes Access to IonQ, OQC, Rigetti; PennyLane integration; QFL possible via custom FL orchestration.
@@ VI-A1 QFL for Cross-Silo Network Security
Cross-Silo QFL lets various organizations (e.g., banks, government agencies or ISPs) collaboratively detect intrusions or attacks without sharing raw network traffic. Each organization has substantial infrastructure (HPC clusters, quantum simulators, or priority access to cloud QPUs). Each silo collects network traffic from its own perimeter. These classical network traffic data are mapped to quantum states using feature mapping techniques, and a quantum model is trained locally across each organization. Only quantum circuit parameters or model weights are securely aggregated at a central coordinator. Techniques such as QKD [scarani2009security] and quantum secure multi-party computation [crepeau2002secure] can be incorporated to ensure the security of the shared-parameters while communicating with the server.
@@ VI-B Communication in QFL
Communication in QFL can be purely classical (with quantum local training), fully quantum (with quantum state sharing), or hybrid. The choice depends on hardware constraints, desired security, and application domain. In the near term, classical communication secured with QKD [mao2018integrating, qi2016simultaneous] is most practical, whereas entanglement-assisted quantum communication represents the long-term vision.
@@
To explore these modalities in more depth, we distinguish between classical communication and quantum communication in QFL [nguyen2025quantum].
@@ VI-C Secure Parameter Sharing in QFL
An efficient and secure communication is one of the major challenges of FL. Classical encryption techniques provide a certain level of protection, but they may fail with increasingly sophisticated attacks. Quantum-secure multi-party computation (QSMPC) [crepeau2002secure] approaches in QFL play a foundational role by enabling multiple parties to jointly compute model updates without revealing their private data. Building upon this, protocols such as Quantum Key Distribution (QKD) [scarani2009security] and Quantum Secure Direct Communication (QSDC) [zhang2017quantum] leverage quantum mechanics principles, like the irreversibility of measurement and photon polarization, to safeguard against eavesdropping by detecting interception through elevated quantum bit error rates. Additionally, quantum teleportation [bouwmeester1997experimental] allows the transfer of quantum states without physically transmitting the qubits themselves. Using pre-shared entanglement combined with classical communication, teleportation facilitates highly secure quantum parameter sharing, especially over long or noisy channels. Finally, quantum techniques such as quantum error correction (QEC) [terhal2015quantum] can be utilized to preserve the integrity and reliability of quantum data exchanged via teleportation, QKD, or QSDC. Various quantum techniques can be used to compress the model to reduce the communication overhead.
@@ VI-D1 FedQNN
The FedQNN framework proposed by Innan et al. in [innan2024fedqnn] extends federated learning by integrating quantum neural networks (QNNs) to enable collaborative model training while maintaining local data privacy. In this approach, each client retains its private dataset and encodes it into quantum states processed through PQCs. The quantum model updates, represented as optimized quantum parameters, are communicated to a central server rather than the raw data. The server aggregates these updates to form a global quantum model, mathematically expressed as the average of client parameters, which enhances overall predictive performance and accuracy. Within the QNN, Hadamard gates create superposition, rotation gates introduce phase shifts and superpositions along X and Y axes, and the $U(\theta)$blocks iteratively optimize network parameters. Measurements on the final quantum state produce predictions, which are evaluated using a mean squared error (MSE) loss function. FedQNN also incorporates secure communication protocols and privacy-preserving aggregation mechanisms, allowing clients to contribute to the global model without exposing sensitive data. This integration of QNNs with federated learning leverages the computational advantages of quantum circuits while addressing the challenges of distributed and privacy-sensitive machine learning environments.
@@ VI-D2 Noise Aware Clustered Quantum Federated Learning (NAC-QFL)
In [sahu2024nac], Sahu et al. proposed NAC-QFL. The NAC-QFL system integrates a QML model with a parameterized quantum circuit into a federated learning framework across a set of quantum devices organized into clusters to optimize communication efficiency. Each cluster has a cluster head responsible for device selection, local model training, and coordination with a central aggregation server. Clustering uses a K-means algorithm with a distance metric combining inter-device communication distance and entanglement-assisted channel capacity, and cluster heads are selected based on both channel capacity and classical computational resources. The selection of devices within clusters is sensitive to noise, and the effective noise of each device is modeled considering qubit relaxation, dephasing, gate errors, measurement errors, and state preparation errors. The least noisy devices that satisfy circuit capacity and parallelization constraints are selected for training. During NAC-QFL training, the global model is distributed from the server to cluster heads, propagated to selected devices for local training, and aggregated back at the server, ensuring scalable, robust, and noise-resilient federated quantum learning.
@@ VI-D3 Multimodal Quantum Federated Learning Framework with Fully Homomorphic Encryption (MQFL-FHE)
Dutta et al. in [dutta2024mqfl] proposed MQFL-FHE, which enables secure federated learning across heterogeneous data modalities, such as images, text, or biomedical sequences, while preserving client privacy. In this framework, each client preprocesses its local multimodal data and encodes it into quantum states using parameterized unitary operations. Local quantum model updates are computed via PQCs and subsequently encrypted using the CKKS fully homomorphic encryption (FHE) scheme, allowing the central server to aggregate encrypted model updates without accessing the raw data. Feature-level fusion is performed locally using attention-based mechanisms to capture correlations across modalities, and the aggregated global model is updated homomorphically by weighting client contributions according to their dataset size. This integration of quantum computation and FHE mitigates the performance degradation typically introduced by encryption while ensuring that sensitive data remains private during the federated learning process. The approach is scalable to multiple modalities and allows each client to contribute securely to a global model that captures multimodal relationships effectively.
@@ VI-D5 Chained Continuous Quantum Federated Learning (ccQFL)
In [gurung2025chained], Gurung et al. proposed ccQFL. In QFL, a central server is responsible for performing aggregation of parameters to produce a global model making it a single point of failure if the server is compromised. The constant communication between a central server and clients may not be feasible in the real world. ccQFL addresses this issue by performing serverless aggregation. Authors have introduced two variants of ccQFL. In the first variant, the devices are randomly grouped or selected based on some selection criteria. The devices train the models in parallel within the group. After the parallel training within the group is completed, the intra-group training is conducted sequentially without the need for an aggregation mechanism. In the second variant, $n$ devices are grouped in pairs. The model is first trained on the first device of each pair and passed on to the second device of the same pair for further training. The trained model from the second device is passed to the untrained second device in alternate pairs. Finally, the models of these devices are averaged and the process continues until the training is completed. This method ensures serverless aggregation of the model and also helps in dealing with heterogeneity and non-IID data.